Two stations A and B are 10 km apart in a straight track, and a train starts from A and

comes to rest at B. For three quarters of the distance, the train is uniformly accelerated and

for the remainder distance uniformly retarded. If it takes 15 minutes over the whole

journey, find its acceleration, retardation and the maximum speed it attains.


1
Expert's answer
2021-05-05T07:45:39-0400


[S1 = 3/4 x 10,000 = 7500 m]



Also, vmax = 22.22 = a1t1

a1 = 22.22 / 675 = 0.0329 m/sec2. (Ans.)

and a2 = 3a1 = 3 x 0.0329 = 0.0987 m/sec2. (Ans.)